Lars:

The equation y = 12x describes the amount of money Louis earns, where x is the number of hours he works, and y is the amount of money he earns. The table shows the amount of money Carl earns for different numbers of hours worked. Carl’s Earnings Time (h) 3, 5, 8, 10 Money earned ($) 45, 75, 120, 150 Create an equation for the amount of money Carl earns.

Arieonna:

you would have to divide $75 ÷ 5 = $15 per hour.

surjithayer:

let the equation be y=ah+b when h=3,y=45 45 =3a+b (1) when h=5,y=75 75=5a+b (2) (2)-(1) gives 2a=30 a=30/2=15 from (1) 45=3*15+b b=45-45=0 y=15h